OpenStudy (anonymous):

Solve: 2 over 7 = 6 over x Solve: 2 over 7 = 6 over x @Mathematics

OpenStudy (anonymous):

x = 21

OpenStudy (anonymous):

\[2/7 = 6/x\] cross multiply \[2x = 42\] divide by 2 \[x = 21\]

OpenStudy (anonymous):

21